OpenStudy (anonymous):

Simplify f(x) = 3x^2-4x+1/3x-1. Using complete sentences, explain why f(1) = 0, f(0) = -1, and f(-1) = -2, yet f(1/3) is undefined. Make sure to show your work.

OpenStudy (anonymous):

if \(x=\frac{1}{3}\) then the denominator will be \(3\times \frac{1}{3}-1=0\) and you cannot divide by zero

OpenStudy (anonymous):

i tried it once, almost didn't survive